OpenStudy (anonymous):

Give the equation of a line that is parallel to the line –3x + 4y = –2 and passes through the point (–4, 6).

OpenStudy (anonymous):

A. y=-3/4x+3 B. y= 3/4x+9 C. y= 3/4x - 1/2 D. y= 4/3x -4

hartnn (hartnn):

first find the slope of line –3x + 4y = –2 can you ?

OpenStudy (anonymous):

Hey ... two lines are parallel if they have the same slopes... So rearrange the given line to get y=.......... The coefficient of 'x' in such case would be the slope of the given line... as y=mx +c has slope 'm'...

OpenStudy (anonymous):

@hartnn m=3/4

hartnn (hartnn):

thats is correct! now since parallel lines have same slope, slope of your line also = m=3/4 you have point (x1,y1) = (-4,6) use the form y-y1 = m(x-x1) to get the equation....

OpenStudy (anonymous):

@hartnn y=3/4x + 9

OpenStudy (anonymous):

is this the correct answer

hartnn (hartnn):

yes, it is :)

OpenStudy (anonymous):

Thank you!! :)

hartnn (hartnn):

Welcome!! :)

OpenStudy (anonymous):

Parallel lines have the same gradient or slope so by re-arrenging your eqn...we get y=3x/4-1/2 which is in the form y=mx+c Hence 3/4 is the slope.... By using the formular for slope....m=(Y-y1)/(X-x1) where x1=-4 and y1=6 we subsitute into the formular....3/4=(y-6)/(x-_4) by working that out we get....y=3x/4+9